OpenStudy (anonymous):

If James received a 9% raise and is now making $102,000 per year, what was his salary before the raise? Round your answer to the nearest dollar.

OpenStudy (anonymous):

x = 102000 - 0.09x 1.09x = 102000 x = 102000/1.09 x = 93578